### v1

# Unified Source-Free Domain Adaptation

Abstract
In the pursuit of transferring a source model to a target domain without access to the source training data, Source-Free Domain Adaptation (SFDA) has been extensively explored across various scenarios, including closed-set, open-set, partial-set, and generalized settings. Existing methods, focusing on specific scenarios, not only address only a subset of challenges but also necessitate prior knowledge of the target domain, significantly limiting their practical utility and deployability. In light of these considerations, we introduce a more practical yet challenging problem, termed unified SFDA, which comprehensively incorporates all specific scenarios in a unified manner. To tackle this unified SFDA problem, we propose a novel approach called Latent Causal Factors Discovery (LCFD). In contrast to previous alternatives that emphasize learning the statistical description of reality, we formulate LCFD from a causality perspective. The objective is to uncover the causal relationships between latent variables and model decisions, enhancing the reliability and robustness of the learned model against domain shifts. To integrate extensive world knowledge, we leverage a pre-trained vision-language model such as CLIP. This aids in the formation and discovery of latent causal factors in the absence of supervision in the variation of distribution and semantics, coupled with a newly designed information bottleneck with theoretical guarantees. Extensive experiments demonstrate that LCFD can achieve new state-of-the-art results in distinct SFDA settings, as well as source-free out-of-distribution generalization. Our code and data are available at https://github.com/tntek/source-free-domain-adaptation.
Index Terms:
Unified domain adaptation, source-free, statistical association, latent causal factors, information bottleneck.
1 Introduction
Due to escalating demands for privacy and information protection in public security and commercial competition, well-annotated training data are recognized as crucial for varying purposes, requiring stringent access control. However, in practical scenarios, the use of a source model (pre-trained on the source domain) without access to the actual source data, referred to as Source-Free Domain Adaptation (SFDA) [1], has become more feasible. This approach involves the challenges posed by strict data access controls and emphasizes the adaptation of a pre-trained model to new domains without relying on the availability of the original source data.
SFDA has been explored across various scenarios, as summarized in Table I. The most constrained scenario, termed closed-set, assumes identical categories of interest in both the source and target domains, focusing on addressing covariate shift, such as domain-specific appearance distributions [2]. Generalized SFDA extends this by requiring the adapted model to perform well in both the source and target domains without forgetting the source domain [3]. Moving beyond the vanilla category identity constraint, open-set SFDA [4] considers the target domain with additional categories, while partial-set SFDA [5] takes the opposite direction. Both scenarios require handling the additional challenge of semantic shift.
Typically, prevalent SFDA methods tend to concentrate on particular scenarios, tackling only a subset of challenges. For instance, SFDA-DE[6] focuses on closed-set scenarios, SF-PGL[7] on open-set scenarios, PSAT[8] on generalized settings, and CRS [9] on partial-set scenarios. Nevertheless, this approach significantly restricts their practical utility and deployability, given that we often possess limited prior knowledge about the target domain and minimal control or selection over its conditions.
TABLE I: Summary of different SFDA settings. $C_{s}$/$C_{t}$: The category set of the source/target domain.
Setting Covariate shift Semantic shift Anti-forgetting Category
Closed-set $C_{s}=C_{t}$
Generalized $C_{s}=C_{t}$
Open-set $C_{s}\subset C_{t}$
Partial-set $C_{s}\supset C_{t}$
Unified (Ours) Any
To address the aforementioned limitation, this study introduces a more realistic a more realistic yet challenging problem, referred to as Unified Source-Free Domain Adaptation (Unified SFDA), with the goal of comprehensively addressing all the specific scenarios mentioned earlier in a unified manner. To achieve this, we propose a novel approach called Latent Causal Factors Discovery (LCFD), going beyond the conventional statistical association learning of related variables by exploring the underlying causal relationships [10]. The discovered causal mechanism is expected to be more reliable under varying distributions and semantic contexts, providing a unified solution for distinct SFDA scenarios.
In the absence of label supervision for both distributional and semantic variations, our LCFD is specifically formulated using a structural causal model in the logits space. The latent causal factors are disentangled into two complementary parts: (i) external causal factors and (ii) internal causal factors, consider that the information of both domain is not necessarily complete. For the former, we leverage a pre-trained large Vision-Language (ViL) model with rich knowledge such as CLIP [11], which has been exposed to a vast amount of multimodality information sources. The latter is identified under the guidance of discovered external causal factors. The discovery of external and internal causal factors is alternated by designing a self-supervised information bottleneck with theoretical guarantees.
Our contributions are summarized as follows:
  • (i)
    For the first time, we introduce a unified SFDA problem that comprehensively incorporates various specific scenarios. This design eliminates the need for prior knowledge about the target domain before model deployment, enhancing practical usability and applicability in real-world scenarios.
  • (ii)
    We introduce a novel approach for unified SFDA, termed LCFD. Instead of learning the statistical description of the problem reality as conventional methods do, our LCFD is formulated under the causality perspective. It aims to reveal the underlying causal relationships between latent variables and model decisions, providing favorable robustness against both distributional and semantic shifts.
  • (iii)
    Extensive experiments demonstrate that LCFD consistently achieves new state-of-the-art results on varying SFDA scenarios as well as source-free out-of-distribution generalization.
2 Related Work
2.1 Source-Free Domain Adaptation
Most of the prior SFDA methods consider the closed-set setting where the source and target domains share the same classes and the focus is cross-domain distribution alignment. Existing methods exist in two categories. The first converts SFDA to Unsupervised Domain Adaptation (UDA) by constructing a pseudo-source domain [12], exploiting source prototype-guided data generation [13] or splitting source-like subset from target data [14]. The second follows the paradigm of self-supervised learning [15, 16], introducing self-guidance. Besides widely used pseudo-labels [15, 17], geometry information [18, 19, 20], constractive data [9] and historical data [8] have also been exploited.
Recently, more practical settings have been studied. For example, the generalized setting shifts aims to mitigate forgetting the source domain [3]. A representative approach is combining continue learning techniques with cross-domain adaptation, e.g., the domain attention-based gradient regulation [3] and source guidance constructed by historical information refining [8]. There exist works considering open-set [7] and partial-set [21] settings. For instance, to control the open-set risk, a progressive balanced pseudo-labeling strategy [7] and closed-set class prototypes are proposed [22]. Estimating the target data distribution helps with negative transfer reduction from the partial classes shift [23].
Alternatively, some works were proposed to enhance the discriminating ability for out-of-distribution classes [24]. There are two main strategies: Expanding inter-class distance to prevent semantic confusion [25, 15], and introducing external semantics, e.g., CLIP, to match out-of-distribution [26].
Existing works as above focus on a specific SFDA setting, limiting their usability and generality substantially in practice. To address this limitation, we introduce a unified SFDA setting that incorporates all the pre-existing scenarios. With this new setup, we aim to foster more advanced SFDA methods that could generalize to a variety of problem settings without the need for designing and maintaining an array of distinct algorithms.
2.2 Causality methods in transfer learning
Causality methods aim for a robust function relation between random variables, which is insensitive to the extra variation/disturbing [27]. This property leads to recent popularity of this new paradigm in transfer learning, e.g., UDA [28], domain generalization (DG) [29] and out of distribution (OOD) [30]. These attempts aim to reduce non-causal factors by introducing artificial intervention. For example, the probability distribution change (used to represent the domain shift) and semantic consistency constraint are jointly employed to implement this intervention. As source labels are available to ensure consistency, these methods put the highlight on constructing the distribution change, following in two lines. One explicitly generates augmented data to disturb the distribution, e.g., the non-linear augmentation and spatially-variable blending augmentation [31], inverse Fourier transformation augmentation [32], and generative models-based cross-domain image style transformation [33, 34]. The other implicitly modifies the distribution by exploiting cross-domain data. For example, MatchDG [35] developed cross-domain contrastive learning where input’s positive sample was randomly selected from samples with the same class. COR [30] adopted the variational inference encoder to infer unobserved causal factors from historical data, taking the time domain as a natural distribution change.
Compared with the existing works, the following three features distinguish LCFD from them. (i) LCFD does not rely on real labels that are indispensable for previous works. (ii) LCFD discovers the causal factors instead of indirectly removing the non-causal ones. (iii) LCFD builds the structural causal model at the logits level whilst previous methods construct on raw data.
Figure 1: Comparison of Structural Causal Models (SCM). (a) SCM for transfer learning with source domain, where $S$ and $U$ represent causal and non-causal factors. Domain shift is generally caused by $U$, which, together with the generalizable information $S$, e.g., the shape/structure of a dog (see Fig. 2), form the observation $x$. (b) Our proposed SCM for SFDA in a latent space (e.g., the logit $\boldsymbol{a}$) where the causal factors $S$ are decomposed into the external $S_{e}$ and internal $S_{i}$ components.
3 Methodology
3.1 Unified SFDA problem
SFDA aims to transfer a model pre-trained on the source domain to a different but related target domain without labelling. Formally, let $\mathcal{X}_{s}\!=\!\{\boldsymbol{x}_{i}^{s}\}_{i=1}^{n_{s}}$ and $\mathcal{Y}_{s}\!=\!\{{y}_{i}^{s}\}_{i=1}^{n_{s}}$ be the source samples and their truth labels. Similarly, the unlabeled samples and the truth target labels are $\mathcal{X}\!=\!\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ and $\mathcal{Y}\!=\!\{{y}_{i}\}_{i=1}^{n}$, respectively. The unified SFDA task is to learn a target model $f_{\theta_{t}}:\mathcal{X}\to\mathcal{Y}$ without any prior knowledge of the target domain not the relationships between the source and target domains (see Tab. I), given (i) the source model $f_{\theta_{s}}$ pre-trained over $\mathcal{X}_{s}$, $\mathcal{Y}_{s}$, and (ii) the unlabeled target data $\mathcal{X}$.
3.2 Causality in transfer learning with source domain
In transfer learning with source domain available, such as DG, and UDA, the causality relation hidden in the statistical dependence (between raw image $\boldsymbol{x}$ and available label $\boldsymbol{y}$) can be summarized to a Structural Causal Model (SCM). As depicted in Fig. 1(a), $\boldsymbol{x}$ is generated jointly by non-causal factors ${U}$ and causal factors ${S}$, while label $\boldsymbol{y}$ is determined by ${S}$ alone. Here, non-causal factors collectively represent the latent variables that generate the category-independent appearance of images, e.g., different domain styles and spurious dependence (Fig. 2); The causal factors present the ones determining the classification, e.g., the object shape.
In a causal view, solving the transfer learning involves eliminating the domain shift by building a robust classification function, i.e., $P(\boldsymbol{y}|S)$. With SCM, we need to disconnect $U$ to $\boldsymbol{x}$. So existing works impose intervention upon $\boldsymbol{x}$, according to the intervention theory [10]. Formally, this scheme can be formulated as
$\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|do(x),S),\end{split}$ (1)
where $do(\cdot)$ means the do-operation of imposing intervention upon variables.
In summary, the key idea is when we change the probability distribution of non-causal factors (intervention) whilst keep the category (prediction consistency), the causality can be extracted. Often, domain related data augmentation is adopted as intervention, whilst the source labels ensure prediction consistency. However both conditions are unavailable in SFDA, making it inapplicable.
3.3 Latent Causality in Unified SFDA
To address the challenges outlined earlier, we propose an approach named Latent Causal Factors Discovery (LCFD), taking into account the following considerations. Instead of extracting causality in the raw input space $\boldsymbol{x}$, we shift our focus to the logit space ${\mathcal{A}}$: $\boldsymbol{a}=f_{\theta_{t}}(\boldsymbol{x})\in\mathcal{A}$. The choice of this space for causality analysis is crucial for two main reasons. First, at the presence of domain distribution gap during adaptation, using the logits essentially imposes a probability distribution perturbation, which could be considered as a special sort of intervention. This post-intervention property facilitates the extraction of causal factors [35, 30, 10]. Second, integration of additional information sources becomes simpler in this semantic space, without being constrained by distinct model architectures.
Figure 2: Illustration of domain shift. (a) Different appearance styles of dog; (b) Spurious dependence of background: Cow in grass ground vs. on a car.
Figure 3: Overview of our proposed LCFD framework: (a) Phase-1: Discovering the external causal factors $S_{e}$ in form of prompt context $\boldsymbol{v}_{e}$ from a frozen ViL model using our self-supervised information bottleneck algorithm; (b) Phase-2: Discovering the internal causal factors $S_{i}$ where the updated prompt context $\boldsymbol{v}_{e}$ is used to predict pseudo-labels as the prior information.
An overview of our LCFD is depicted in Fig. 1(b). It models non-causal $U$ and causal $S$ factors in a latent space ${\mathcal{A}}$. Uniquely, the causal part is further decomposed into internal $S_{i}$ and external $S_{e}$ components, consider that the information of source and target domain is not necessarily comprehensive. In contrast to conventional intervention driven learning, we focus on discovering $S$ from the intertwined $S$ and $U$ as
$\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|dis(S|\{U,S\})),~{}S=\{S_{i},S_{e}\}.\end{split}$ (2)
To make this process computationally tractable, we consider
Principle 1
Independent Causal Mechanisms Principle [36]: The conditional distribution of each variable given its causes (i.e., its mechanism) does not inform or influence the other mechanisms.
This principle means $S_{i}$ and $S_{e}$ are independent in our context. So we can simplify our above formula as:
$\small\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|dis(S_{e}|\{U,S_{e}\}))\cdot P(\boldsymbol{y}|dis(S_{i}|\{U,S_{i}\})).\end{split}$ (3)
3.4 LCFD Design
We present a concrete design for realizing the LCFD as formulated above. An overview is depicted in Fig. 3. By Eq. (3), we implement the causality discovery in two parts: phase-1 for $S_{e}$ and phase-2 for $S_{i}$, as detailed below.
3.4.1 External Causality Discovery
To obtain external causality for a target task, we explore the potential of recent large multimodal ViL models such as CLIP [11]. This is because they have experienced a vast amount of knowledge in pretraining, and exhibited general ability in many different perception tasks.
Specifically, we accomplish this idea by encoding $S_{e}$ into the prompt context $\boldsymbol{v}_{e}$ of a ViL model $g_{\theta}$ efficiently. That is, our learning target is $\boldsymbol{v}_{e}$, which is the explicit expression of $S_{e}$. We start by formulating the $S_{e}$ discovery as maximizing the correlation between a random variable representing $S_{e}$ and the prediction of the target model. Then, the original formulation is converted to a self-supervised information bottleneck problem with theoretical guarantee, followed by the deep learning instantiation.
With the target samples $\mathcal{X}$ and learnable $\boldsymbol{v}_{e}$, we obtain the prediction of $g_{\theta}$ as $\hat{\mathcal{P}}=\{\hat{\boldsymbol{p}}_{i}\}_{i=1}^{n}$. Meanwhile, we obtain the logits of target model $f_{\theta_{t}}$ for the same data as ${\mathcal{A}}=\{{\boldsymbol{a}}_{i}\}_{i=1}^{n}$. We introduce three random variables, $V$, $Y$, and $Z$, following the probability distributions of $\boldsymbol{v}_{e}$, $\hat{\mathcal{P}}$ and ${\mathcal{A}}$, respectively. In LCFD, we further write $Z=\{Z_{e},Z_{i}\}$ where the latent random variables $Z_{e}$ and $Z_{i}$ represent the external $S_{e}$ and the internal $S_{i}$ causal factors, respectively.
In information-theoretic view, for causality relationships we have to ensure the correlation between ${Z}_{e}$ and $V$. This could be considered an approximation with computational ease. Formally, discovering $S_{e}$ needs maximizing the following mutual information:
$\begin{split}\max_{\boldsymbol{v}_{e}}I\left({{Z_{e}},V}\right),\end{split}$ (4)
where function $I(\cdot,\cdot)$ computes mutual information of two random variables.
Consider that prompt $V$ is specific to the ViL model $g_{\theta}$, we further transform the above formula as
$\begin{split}\max_{\boldsymbol{v}_{e}}I\left({{Z_{e}},Y}\right),Y=\phi\left(g_{\theta}\left(V\right)\right),\end{split}$ (5)
where $\phi(\cdot)$ is softmax function; so that the two variables under correlation maximization reside in similarly semantic spaces. Theoretically, we prove that this transformed is a low bound of the original, thus valid in terms of optimization.
First, we have the following Lemma 6 (see the proof in Appendix).
Lemma 1
Given random variables $Z_{1}$, $X_{1}$ and $Y_{1}$ where $X_{1}$, $Y_{1}$ satisfy a mapping $f_{1}:X_{1}\mapsto Y_{1}$. When $f_{1}$ is compressed, i.e., the output’s dimension is smaller than the input’s,
$\begin{split}I\left({{Z_{1}},X_{1}}\right)\geq I\left({{Z_{1}},Y_{1}}\right).\end{split}$ (6)
The ViL model $g_{\theta}$ is indeed compressing, as it maps a high-dimensional image to a low-dimensional category vector. Letting $Z_{1}=Z_{e}$, $X_{1}=X$, $Y_{1}=Y$ and $f_{1}=\phi\left(g_{\theta}(\cdot)\right)$, we have $I({{Z_{e}},V})\geq I({{Z_{e}},Y})$ according to Lemma 6.
Figure 4: Realising our self-supervised information bottleneck w.r.t. Theorem 8. The item $\rm{T}_{1}$ is estimated by variational information maximization with Gaussian distribution assumption $\mathcal{N}(Z,\Sigma_{Z})$. The item $\rm{T}_{2}$ is computed by mutual information definition where the function $f_{w}$ integrates the min-max optimization, using covariance matrix $\Sigma_{Z}$ from $\rm{T}_{1}$ as the weighting parameters $w$.
Self-supervised information bottleneck. In Eq. (5), $Z_{e}$ is hidden part of the observable logtis along with non-causal and internal causal factors. To facilitate this learning process of $Z_{e}$, we derive a self-supervised information bottleneck algorithm without prior assumption on the distribution [37]. In general, information bottleneck conducts min-max optimization to squeeze out the essentials w.r.t. a task. We introduce a random variable $Z{{}^{\prime}}$ (i.e., the bottleneck) for the ViL’s logits $\widehat{\mathcal{O}}=\{\hat{\boldsymbol{o}}_{i}\}_{i=1}^{n}$ between $V$ and $Y$, thereby forming a bottleneck style expression as:
$\begin{split}\min_{\boldsymbol{v}_{e}}I\left({Z,Z{{}^{\prime}}}\right)-I\left({Z{{}^{\prime}},Y}\right),\end{split}$ (7)
where $Z$ is the logit variable of target model $f_{\theta_{t}}$ which is observed. Note, we do not include the item $I\left({Z{{}^{\prime}},Y}\right)$ in minimization due to lacking of supervision signal like ground-truth labels as used in conventional information bottleneck [38], i.e., a self-supervised formulation. However, we theoretically prove that Eq. (7) is an upper bound as below thus optimizationally valid.
Theorem 1
Suppose that there are five random variables $Z$, $V$, $Z{{}^{\prime}}$, $Y$ and $Y{{}^{\prime}}$. Among them, $Z$ represents the target domain knowledge; $V$, $Z{{}^{\prime}}$ and $Y$ express the input instances, an intermediate features of the ViL model and predictions, respectively; $Y{{}^{\prime}}$ depicts a pseudo-label that has a functional relationship $f_{w}$ ($w$ is learnable parameters) with $Z{{}^{\prime}}$. When $f_{w}$ is reversible and uncompressed, we have this upper bound
$\small\begin{split}&\min_{\boldsymbol{v}_{e}}I\left({Z,{Z{{}^{\prime}}}}\right)-I\left({{Z{{}^{\prime}}},Y}\right)\\ &\leq\min_{\boldsymbol{v}_{e},w}\underbrace{I\left({Z,{Z{{}^{\prime}}}}\right)}_{\rm{T_{1}}}-\underbrace{I\left({{Y{{}^{\prime}}},Y}\right),{Y{{}^{\prime}}}=f_{w}\left({{Z{{}^{\prime}}}}\right)}_{\rm{T_{2}}}.\end{split}$ (8)
The proof is given in Appendix LABEL:sec-thm. Theorem 8 suggests that the desired latent factors refinement is conditionally equivalent to a self-supervised information bottleneck, when pseudo-label $Y{{}^{\prime}}$ is generated from the intermediate ViL feature $Z{{}^{\prime}}$ without information loss.
We realize the proposed information bottleneck using variational information maximization [39]. The key is to construct the function $f_{w}$ with three properties: (i) The output is a probability distribution, e.g., including a softmax operation $\phi$. (ii) No information loss, e.g., no zero weights in $w$ for weighting based functions. (iii) The weighting parameters $w$ are constrained to be relevant to $I(Z{{}^{\prime}},Z)$. The instantiation of $T_{1}$ and $T_{2}$ are elaborated below.
As exactly computing mutual information of two vector variables is intractable, we maximize the variational lower bound for the term $T_{1}$ (see the proof in [40]):
$\small\begin{split}I\left(Z,Z{{}^{\prime}}\right)&=\mathbb{E}_{{Z,{Z{{}^{\prime}}}}}\left[\log\frac{q\!\left({Z{{}^{\prime}}}|Z\right)}{p\left({Z{{}^{\prime}}}\right)}\right]+\displaystyle{\rm{KL}}\left(\left.{q\!\left({Z{{}^{\prime}}}|Z\right)}\right\|p\left({Z{{}^{\prime}}}|Z\right)\right)\\ &\geq H\left({Z{{}^{\prime}}}\right)+\mathbb{E}_{{Z,{Z{{}^{\prime}}}}}\left(\log q\left({Z{{}^{\prime}}}|Z\right)\right),\end{split}$ (9)
where $q({Z{{}^{\prime}}}|Z)$ is variational distribution approximating the real distribution $p({Z{{}^{\prime}}}|Z)$, ${\rm{KL}}(\cdot,\cdot)$ is ${\rm{KL}}$-divergence function, and $H({Z{{}^{\prime}}})$ is a constant. In practice, we model $q({Z{{}^{\prime}}}|Z)$ as a Gaussian distribution $\mathcal{N}(Z,\Sigma_{Z})$ with mean $Z$ and diagonal covariance matrix $\Sigma_{Z}$. Then, the loss for minimizing $T_{1}$ can be formulated (see $\rm{\mathbf{T}}_{1}$ in Fig. 4):
$\begin{split}&\mathcal{L}_{\rm{VMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)\\ &=-\mathbb{E}_{{Z,{Z{{}^{\prime}}}}}\left(||{Z{{}^{\prime}}}-Z||_{\Sigma_{Z}^{-1}}^{2}+\log|\Sigma_{Z}|\right)\\ &=-\frac{1}{n}\sum_{i=1}^{n}\left(\hat{\boldsymbol{o}}_{i}-{\boldsymbol{a}}_{i}\right)^{T}{\Sigma_{Z}^{-1}}\left(\hat{\boldsymbol{o}}_{i}-{\boldsymbol{a}}_{i}\right)+\log{|\Sigma_{Z}^{i}|}.\end{split}$ (10)
Considering that the ViL’s prediction $Y$ and the pseudo-label $Y{{}^{\prime}}$ are probability distributions, we compute the second term $\rm{T_{2}}$ by the original mutual information [41]: $I(Z,Z{{}^{\prime}})={p(Z,Z{{}^{\prime}})}\mathord{\left/{\vphantom{{}{}}}\right.\kern-1.2pt}{[p(Z)p(Z{{}^{\prime}})]}$. The corresponding loss function is
$\begin{split}&\mathcal{L}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)=\frac{1}{n}\sum_{i=1}^{n}{I}\left(\hat{\boldsymbol{p}}_{i},{\boldsymbol{p}}^{\prime}_{i}\right),\\ &\hat{\boldsymbol{p}}_{i}=\phi(\hat{\boldsymbol{o}}_{i}),~{}~{}{\boldsymbol{p}}^{\prime}_{i}=f_{w}(\hat{\boldsymbol{o}}_{i})\end{split}$ (11)
where $f_{w}$ is a weighting-based function we construct as follows.
From Eq. (10), $I(Z{{}^{\prime}},Z)$ is a Mean-Square Error (MSE) weighted by $\Sigma_{Z}$. That means $\Sigma_{Z}$ detects dimension importance duration minimization, with high-weight dimensions to be suppressed, thus corresponding to the non-causal factors $U$ in our formulation. Thus, taking ${{1}\mathord{\left/{\vphantom{{}{}}}\right.\kern-1.2pt}{\Sigma_{Z}}}$ as the weighting parameters of $f_{w}$ shifts the focus of predicting to the causal component $S_{e}$ (see $\rm{\mathbf{f}}$ in Fig. 4), formally,
$\begin{split}{\boldsymbol{p}}^{\prime}_{i}=f_{\left[w=\Sigma_{Z}^{i}\right]}\left(\hat{\boldsymbol{o}}_{i}\right)=\phi\left(\frac{1}{\Sigma_{Z}^{i}}\right)\odot\phi\left(\hat{\boldsymbol{o}}_{i}\right).\end{split}$ (12)
Combining Eq. (10), (11) and (12) together, our self-supervised objective for external causality discovery is formed as
$\begin{split}&\min_{{\boldsymbol{v}_{e}},\Sigma_{Z}}\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)\\ &=\mathcal{{L}}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)-\alpha\mathcal{{L}}_{\rm{VMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right).\end{split}$ (13)
Remark. Our design differs from existing information bottleneck in two aspects: While previous information extraction [42] is confined to a single model, ours is of cross-model. Our formulation is self-supervised, in contrast to previous supervised designs.
Algorithm 1 Training LCFD
Input: Unlabelled target data $\mathcal{X}$, pre-trained source model $f_{\theta_{s}}$, a frozen ViL model $g_{\theta}$, max training iteration number $M$, prompt context $\boldsymbol{v}_{e}$.
1:  Initialize the target model $f_{\theta_{t}}$ by $f_{\theta_{s}}$
2:  Initialize $\boldsymbol{v}_{e}$ with a fixed template
3:  for iter $k=1$ to ${M}$ do
4:     Sample a mini-batch $\mathcal{X}_{b}$ from $\mathcal{X}$.
5:     Discover external causal factors $S_{e}$: update $(\boldsymbol{v}_{e},\Sigma_{Z})$ by optimizing $\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X}_{b};\boldsymbol{v}_{e},\Sigma_{Z}\right)$, fixing $\mathcal{{L}}_{\rm{IC}}$.
6:     Convert updated $\boldsymbol{v}_{e}$ to target pseudo-labels for this batch $\mathcal{X}_{b}$, i.e., obtain ${\mathcal{Q}}_{b}=\phi(g_{\theta}(\mathcal{X}_{b},\boldsymbol{v}_{e})$.
7:     Discover internal causal factors $S_{i}$: update $f_{\theta_{t}}$ by optimizing $\mathcal{{L}}_{\rm{IC}}$, fixing $\mathcal{{L}}_{\rm{EC}}$.
8:  end for
9:  return: The adapted model $f_{\theta_{t}}$.
3.4.2 Internal Causality Discovery
In phase-2, we subsequently discover the internal causal factors $S_{i}$ by updating the target model. That is, we encode $S_{i}$ into the target model as it is latent in nature. This is achieved using the pseudo-labels predicted by the ViL model and prompt context in a self-supervised learning manner.
Formally, denote ${\mathcal{Q}}=\phi(g_{\theta}(\mathcal{X},\boldsymbol{v}_{e}))$ the pseudo-labels of the target data $\mathcal{X}$ generated by the ViL model $g_{\theta}$, using $\boldsymbol{v}_{e}$ updated in phase-1, and ${\mathcal{P}}=\phi(f_{\theta_{t}}(\mathcal{X}))$ the predictions of $\mathcal{X}$ by the target model $f_{\theta_{t}}$. We design the objective function for discovering $S_{i}$ as
$\footnotesize{{\cal L}_{{\rm{IC}}}}\mathop{=}\min_{{\theta_{t}}}\underbrace{H({\cal P})+\tau\sum\limits_{c=1}^{C}{{{{\rm{KL}}}}}\left({\left.{{\varrho_{c}}}\right\|{1\over C}}\right)}_{{{\cal L}_{{\rm{UN}}}}}+~{}\sigma\underbrace{{{\cal L}_{{\rm{SCE}}}}\left({{\cal P},{\cal Q};{\theta_{t}}}\right)}_{{{\cal L}_{{\mathop{\rm PL}\nolimits}}}}$ (14)
where ${H}({\mathcal{P}})\!=\!-\sum_{i=1}^{n}\boldsymbol{p}_{i}\log\boldsymbol{p}_{i}$ estimates the information entropy, $\frac{1}{C}$ stands for the uniform distribution over $C$ categories, ${\varrho}_{c}=\frac{1}{n}\sum_{i=1}^{n}{p}_{i,c}$ is empirical label distribution, ${p}_{i,c}$ is the probability of $\boldsymbol{x}_{i}$ in the $c$-th category, ${{\cal L}_{{\rm{SCE}}}}$ is the softmax-cross-entropy loss, and $\tau$, $\sigma$ are trade-off parameters. The first term ${\cal L}_{{\rm{UN}}}$ is a regularization for encouraging prediction confidence and avoiding solution collapse [43]. Aligning with the pseudo labels estimated using the prompt context, the second item ${\cal L}_{{\rm{PL}}}$ essentially integrates the external causal factors $S_{e}$ while learning $S_{i}$.
3.5 Model Training
We train LCFD in an alternating manner. In each iteration, we first optimize prompt context $\boldsymbol{v}_{e}$ with ${\cal L}_{{\rm{EC}}}$ whilst the target model is frozen (i.e., phase-1), followed by training the target model by ${\cal L}_{{\rm{IC}}}$ (i.e., phase-2) while the prompt context is frozen. The ViL model is always frozen throughout. More details for training is presented in Alg. 1.
TABLE II: Closed-set SFDA results (%) on Office-Home and VisDA. SF and M means source-free and multimodal, respectively.
Method Venue SF M Office-Home VisDA
[…]
Source 43.7 67.0 73.9 49.9 60.1 62.5 51.7 40.9 72.6 64.2 46.3 78.1 59.2 49.2
CLIP-RN ICML21 51.9 81.5 82.5 72.5 81.5 82.5 72.5 51.9 82.5 72.5 51.9 81.5 72.1 83.7
CLIP-B32 ICML21 59.8 84.3 85.5 74.6 84.3 85.5 74.6 59.8 85.5 74.6 59.8 84.3 76.1 82.9
DAPL-RN TNNLS23 54.1 84.3 84.8 74.4 83.7 85.0 74.5 54.6 84.8 75.2 54.7 83.8 74.5 86.9
PADCLIP-RN ICCV23 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6 88.5
ADCLIP-RN ICCV23W 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9 87.7
SHOT ICML20 55.0 78.7 81.3 69.1 78.9 79.1 68.2 53.6 81.6 73.5 59.4 83.5 71.8 82.9
GKD IROS21 56.5 78.3 82.2 69.2 80.4 78.7 67.4 55.4 82.6 74.3 60.3 84.2 72.5 83.0
NRC NIPS21 57.2 79.3 81.3 68.9 80.6 80.2 66.6 57.3 82.0 71.0 57.9 84.9 72.3 85.9
AaD NIPS22 59.3 79.3 82.1 68.9 79.8 79.5 67.2 57.4 83.1 72.1 58.5 85.4 72.7 88.0
AdaCon CVPR22 47.2 75.1 75.5 60.7 73.3 73.2 60.2 45.2 76.6 65.6 48.3 79.1 65.0 86.8
CoWA ICML22 57.3 79.3 81.0 69.3 77.9 79.6 68.1 56.4 82.6 72.9 61.3 83.7 72.4 86.9
PLUE CVPR23 49.1 73.5 78.2 62.9 73.5 74.5 62.2 48.3 78.6 68.6 51.8 81.5 66.9 88.3
TPDS IJCV23 59.3 80.3 82.1 70.6 79.4 80.9 69.8 56.8 82.1 74.5 61.2 85.3 73.5 87.6
LCFD-C-RN 60.1 85.6 86.2 77.2 86.0 86.3 76.6 61.0 86.5 77.5 61.4 86.2 77.6 89.3
LCFD-C-B32 72.3 89.8 89.9 81.1 90.3 89.5 80.1 71.5 89.8 81.8 72.7 90.4 83.3 89.3
TABLE III: Closed-set SFDA results (%) on DomainNet-126. SF and M means source-free and multimodal, respectively.
[…]
Source 44.6 59.8 47.5 53.3 75.3 46.2 55.3 62.7 46.4 55.1 50.7 59.5 54.7
CLIP-RN ICML21 70.2 87.1 65.4 67.9 87.1 65.4 67.9 70.2 65.4 67.9 70.2 87.1 72.7
CLIP-B32 ICML21 73.5 85.7 71.2 74.7 85.7 71.2 74.7 73.5 71.2 74.7 73.5 85.7 76.3
DAPL-RN TNNLS23 72.4 87.6 65.9 72.7 87.6 65.6 73.2 72.4 66.2 73.8 72.9 87.8 74.8
ADCLIP-RN ICCV23W 71.7 88.1 66.0 73.2 86.9 65.2 73.6 73.0 68.4 72.3 74.2 89.3 75.2
SHOT ICML20 63.5 78.2 59.5 67.9 81.3 61.7 67.7 67.6 57.8 70.2 64.0 78.0 68.1
GKD IROS21 61.4 77.4 60.3 69.6 81.4 63.2 68.3 68.4 59.5 71.5 65.2 77.6 68.7
NRC NIPS21 62.6 77.1 58.3 62.9 81.3 60.7 64.7 69.4 58.7 69.4 65.8 78.7 67.5
AdaCon CVPR22 60.8 74.8 55.9 62.2 78.3 58.2 63.1 68.1 55.6 67.1 66.0 75.4 65.4
CoWA ICML22 64.6 80.6 60.6 66.2 79.8 60.8 69.0 67.2 60.0 69.0 65.8 79.9 68.6
PLUE CVPR23 59.8 74.0 56.0 61.6 78.5 57.9 61.6 65.9 53.8 67.5 64.3 76.0 64.7
TPDS IJCV23 62.9 77.1 59.8 65.6 79.0 61.5 66.4 67.0 58.2 68.6 64.3 75.3 67.1
LCFD-C-RN 75.4 88.2 72.0 75.8 88.3 72.1 76.1 75.6 71.2 77.6 75.9 88.2 78.0
LCFD-C-B32 77.2 88.0 75.2 78.8 88.2 75.8 79.1 77.8 74.9 79.9 77.4 88.0 80.0
4 Experiments
4.1 Implementation Details
Source models. Following [44], for source-free out-of-distribution generalization (SF-OODG) we adopt the Pytorch built-in ResNet50 model (pre-trained on ImageNet) as the source model, taking ImageNet variations as evaluation datasets. For all the other settings, the source model is trained on the source domain in a supervised manner, same as [15, 18, 20].
Networks. Our LCFD involves two networks: the ViL model and the target model. In practice, we choose CLIP [11] as the ViL model. The target model’s structure is the same as the source model. Following [15, 19], a target model consists of a feature extractor and a classifier (a FC layer). We adapt the feature extractor, which pre-trained on ImageNet. For fair comparison, we use ResNet101 as the backbone for VisDA and ResNet50 for the others.
Training. There are three parameters with LCFD: $\alpha$ in $L_{\rm{EC}}$ (Eq. (13)), $\sigma$ and $\tau$ in $L_{\rm{IC}}$ (Eq. (14)). For all settings, we adopt the same configuration $(\alpha,\sigma,\tau)=(0.003,0.4,1.0)$. Here, since the value of $\mathcal{{L}}_{\rm{VMI}}$ is large, $\alpha$ is set to the small one of 0.003. For model training, we use the batch size of 64, SGD optimizer with momentum 0.9 and 15 epochs on all datasets. The prompt template for initialization is ’a photo of a [CLASS].’ [11].
4.2 Closed-Set SFDA
Data sets. In this part, we evaluate LCFD on three challenging benchmarks for domain adaptation. Office-Home [45] is a middle-scale dataset. It contains 15k images belonging to 65 categories from working or family environments, being divided into four domains, i.e., Artistic images, Clip Art, Product images, and Real-word images. VisDA [46] is a large-scale dataset. It includes 12 types of Synthetic to Real transfer recognition tasks. The source domain contains 152k synthetic images, whilst the target domain has 55k real object images from Microsoft COCO. DomainNet-126 [47] is another large-scale dataset. As a subset of DomainNet containing 600k images of 345 classes from 6 domains of different image styles, this dataset has 145k images from 126 classes, sampled from 4 domains, Clipart, Painting, Real, Sketch, as [47] identifies severe noisy labels in the dataset.
TABLE IV: One step adapting results (%) on Office-Home in Generalized SFDA setting. S, T, H are the results on the source and target domains, and harmonic mean accuracy, respectively; the bracket values in red are the gap between S and T; WAD means With Anti-forgetting Design.
Method Venue WAD Avg.
S T (S-T gap) H
Source 98.1 59.2 (38.9) 73.1
SHOT ICML20 84.2 71.8 (12.4) 77.1
GKD IROS21 86.8 72.5 (14.3) 78.7
NRC NIPS21 91.3 72.3 (19.0) 80.4
AdaCon CVPR22 88.2 65.0 (23.2) 74.4
CoWA ICML22 91.8 72.4 (19.4) 80.6
PLUE CVPR23 96.3 66.9 (29.4) 78.4
TPDS IJCV23 83.8 73.5 (10.3) 78.0
GDA ICCV21 80.0 70.2 (9.8) 74.4
PSAT-ViT TMM23 86.4 83.6 (2.8) 85.0
LCFD-C-RN 85.0 77.6 (7.4) 80.7
LCFD-C-B32 86.3 83.3 (3.0) 84.5
TABLE V: Continual adaptation results (%) on Office-Home under Generalized SFDA setting, the first column of each sub-table indicates the adaptation sequence. $\downarrow$ means the average accuracy drop of a test domain on the adaptation path compared with the performance when the domain is first seen.
ResNet50 (RN)
Test Test Test Test
Ar Cl Pr Rw Cl Ar Pr Rw Pr Ar Cl Rw Rw Ar Cl Pr
Ar 97.8 46.1 68.4 71.1 Cl 97.5 52.5 65.8 61.5 Pr 99.6 52.2 40.8 71.3 Rw 98.1 63.1 46.3 78.5
Cl 80.0 63.5 64.3 66.0 Ar 80.1 77.8 70.7 75.6 Ar 91.2 78.4 50.2 78.3 Ar 93.4 78.4 50.4 75.6
Pr 78.4 61.7 85.0 76.2 Pr 75.4 74.4 86.9 78.1 Cl 81.5 71.9 61.5 69.9 Cl 83.0 74.7 62.1 69.7
Rw 83.1 60.2 83.2 84.4 Rw 67.6 78.4 85.4 85.6 Rw 87.3 76.9 59.2 85.6 Pr 87.3 72.5 61.5 85.2
$\downarrow$ 17.3 2.5 1.8 $\downarrow$ 23.1 1.4 1.6 $\downarrow$ 13.0 4.1 2.3 $\downarrow$ 10.2 4.8 0.6
ViT-B/32 (B32)
Test Test Test Test
Ar Cl Pr Rw Cl Ar Pr Rw Pr Ar Cl Rw Rw Ar Cl Pr
Ar 97.8 46.1 68.4 71.1 Cl 97.5 52.5 65.8 61.5 Pr 99.6 52.2 40.8 71.3 Rw 98.1 63.1 46.3 78.5
Cl 84.1 77.0 69.3 72.5 Ar 80.9 83.4 72.3 77.0 Ar 90.0 82.5 49.4 80.1 Ar 94.0 82.8 51.2 76.8
Pr 82.5 70.9 91.0 81.1 Pr 74.2 78.1 90.4 82.2 Cl 83.6 75.6 76.2 76.2 Cl 85.2 73.8 76.2 72.5
Rw 86.3 67.6 88.3 87.7 Rw 69.5 81.3 88.9 89.1 Rw 89.8 79.1 69.5 87.7 Pr 89.5 76.9 71.5 90.0
$\downarrow$ 13.5 7.7 2.7 $\downarrow$ 22.6 3.7 1.5 $\downarrow$ 11.8 5.2 6.7 $\downarrow$ 8.5 7.5 4.7
Competitors. To evaluate LCFD, we select 14 state-of-the-art comparisons divided into three groups. (i) The first is the source model and CLIP that lay the basis for LCFD. (ii) The second is the UDA method DAPL [48], PADCLIP [49], ADCLIP [50] that introduces prompt learning to boost the cross-domain transfer. (iii) The third includes 9 SFDA methods: SHOT [15], GKD [25], NRC [18], SFDE [6], AdaCon [51], CoWA [23], PLUE [17], TPDS [52]. According to idea, SHOT, GKD, CoWA are pseudo-labels-based methods; NRC, PLUE, AdaCon, SFDE exploit the data geometry, e.g., the nearest neighbor, source prototype; TPDS predicts the target probability distribution using error control.
For comprehensive comparisons, we implement LCFD in two variants: (i) LCFD-C-RN (base) and (ii) LCFD-C-B32 (premium). The key distinction lies in the backbone of the CLIP image-encoder. Specifically, for LCFD-C-RN, ResNet101 is employed on the VisDA dataset, while ResNet50 is used on the Office-Home and DomainNet-126 datasets. On the other hand, LCFD-C-B32 adopts ViT-B/32 [53] as the backbone across all datasets. Similarly, all comparison methods using CLIP are marked with the same meaning for a convenience observation.
[…]
On the other hand, LCFD-C-RN defeats CLIP-RN on all tasks, achieving the improvement of 4.5%, 5.6% and 5.3% on Office-Home, VisDA and DomainNet-126 in average, respectively. Similarly, compared with CLIP-B32, the corresponding improvement of LCFD-C-B32 are changed to 7.2%, 6.4% and 3.7%. This indicates the effect of the prompt learning to discover the external causal elements.
4.3 Generalized SFDA
Data sets. This setting evaluates LCFD on the Office-Home, following previous generalized SFDA methods [3, 8].
Evaluation protocol. Unlike closed-set SFDA, the generalized SFDA problem highlights the anti-forgetting ability on the seen source domain. In terms of evaluation rule, the same as [3], we adopt the harmonic mean accuracy that is computed by ${H}=({2*A_{s}*A_{t}})/({A_{s}+A_{t}})$ where $A_{s}$ and $A_{t}$ are the accuracy on the source domain and the target domain, respectively. Note that the $A_{s}$ is computed based on the source-testing set. The same to [3, 8], on the source domain, the ratio of training and testing sets is 9:1.
Competitors. As listed in Tab. IV, two generalized SFDA methods GDA [3], PSAT [8] are additionally selected as the comparisons. In GDA, domain attention was integrated with a local clustering-based self-supervised learning method. Essentially, the capacity for the old tasks is kept by encouraging the model parameters closed to the original values. Unlike GDA, PSAT enforces the adaptation process to remember the source domain by imposing source guidance, building a target domain-centric anti-forgetting mechanism.
One step adapting comparison. Tab. IV reports the results on Office-Home, including H-accuracy, S-accuracy and T-accuracy. In the average H-accuracy aspect, LCFD-C-B32 is only defeated by generalized SFDA method PSAT (having particular design of anti-forgetting mechanism and based on ViT backbone) with a gap of 0.5%. Adopting the weak version LCFD-C-RN still outperforms other methods except for the best PSAT. Although T-accuracy of NRC, CoWA closes LCFD-C-RN due to the high S-accuracy (91%), the S-T gap of them (19%) is much larger than LCFD-C-RN (7.4%). These results imply a better balance of LCFD as one-step adapting.
[…]
4.4 Open-Set & Patial-Set SFDA
Data sets. In this setting, we adopt Office-Home as the evaluation benchmark, as done in [15, 9].
TABLE VI: Partial-set SFDA and open-set SFDA results on the Office-Home dataset (%).
Partial-set SFDA Open-set SFDA
Method Venue Avg. Method Venue Avg.
Source 62.8 Source 46.6
SHOT ICML20 79.3 SHOT ICML20 72.8
HCL NIPS21 79.6 HCL NIPS21 72.6
CoWA ICML22 83.2 CoWA ICML22 73.2
AaD NIPS22 79.7 AaD NIPS22 71.8
CRS CVPR23 80.6 CRS CVPR23 73.2
LCFD-C-RN 82.9 LCFD-C-RN 79.6
LCFD-C-B32 85.8 LCFD-C-B32 83.4
Competitors. Besides methods Source, CLIP, SHOT, CoWA from the closed-set setting, this part adds three new comparisons HCL [54], AaD [55] and CRS [9] that all follow the contrastive learning framework.
Comparison results. As the comparison results on Office-Home presented in Tab. VI, in average accuracy, LCFD-C-B32 surpasses not only all comparisons with an improvement of 2.6% and 10.2% compared with previous work CoWA (partial-set) and CRS (open-set). As for LCFD-C-RN, only CoWA beat it in the partial-set setting with a tiny leading of 0.3%. The quantitative results indicate that LCFD is still competitive on the partial-set and open-set settings.
4.5 Source-Free Out-Of-Distribution Generalization
Data sets. This part evaluates LCFD in the SF-OODG setting based on four challenging ImageNet Variants dataset published in recent years. Specifically, IN-V2 (i.e., ImageNet-V2) [56] is an independent set consisting of natural images, collected from different sources, totally containing 10k images of 1000 ImageNet categories. IN-A (i.e., ImageNet-A) [57] is a challenging set consisting of these ”natural adversarial examples” (misclassified by a vanilla ResNet50 [58]), containing 7.5k images of 200 ImageNet categories. IN-R (i.e., ImageNet-R) [59] is an ImageNet sub-set with artistic renditions. It contains 30k images covering 200 ImageNet categories. IN-K (i.e., ImageNet-Sketch) [60] collects black and white sketches from the ImageNet validation set. This dataset contains 50k images from 1000 ImageNet categories.
Competitors. As shown in Tab. VII, besides these comparison methods adopted in the closed-set SFDA setting, we add four new comparisons, CoOP [44], CoCoOp [61], TPT [26] and ProGrad [62]. Among them, CoOP, CoCoOp, ProGrad perform prompt learning under the supervision of a few labels, whilst TPT proposes self-supervised prompt tuning via minimizing the multi-view entropy. Like the closed-set setting, the CLIP-based methods are marked with the backbone of image-encoder in CLIP.
Comparison results. Tab. VII reports the SF-OODG results on ImageNet Variations. In terms of average accuracy, our methods, LCFD-C-NR and LCFD-C-B32, are significantly ahead of the previous SFDA methods, improving by 10% at least. For CLIP and the CLIP-based methods, the advantage of LCFD is still noticeable. Taking results of the weak version as an example, while improving by 7.0% compared with CLIP-C-RN, LCFD-C-NR surpasses the second-best CLIP-based method TPT-RN by 3.8%.
TABLE VII: SF-OODG results (%) on ImageNet variants. L and M means label-required and multimodal, respectively.
[…]
IN-V2 IN-K IN-A IN-R Avg.
Source 62.7 22.2 0.7 35.1 30.2
CLIP-RN ICML21 51.5 33.3 21.8 56.1 40.7
CLIP-B32 ICML21 54.8 40.8 29.5 66.2 47.8
CoOP-RN IJCV21 55.4 34.7 23.1 56.6 42.4
CoOP-B32 IJCV21 58.2 41.5 31.3 65.8 49.2
CoCoOp-RN CVPR22 55.7 34.5 23.3 57.7 42.8
CoCoOp-B32 CVPR22 56.6 40.7 30.3 64.1 47.9
TPT-RN NIPS22 54.7 35.1 26.7 59.1 43.9
ProGrad-RN ICCV23 54.7 34.4 23.1 56.8 42.2
ProGrad-B32 ICCV23 57.4 41.7 31.9 66.5 49.4
SHOT ICML20 62.5 38.4 2.1 42.7 36.4
GKD IROS21 62.3 38.2 2.1 42.2 36.2
NRC NIPS21 61.5 0.9 1.3 28.9 23.1
AdaCon CVPR22 50.9 18.5 2.0 38.6 27.5
CoWA ICML22 62.7 37.8 1.7 46.3 37.1
PLUE CVPR23 53.0 18.5 2.0 38.6 28.0
TPDS IJCV23 62.9 35.0 2.9 48.4 37.3
LCFD-C-RN 64.4 42.6 22.3 61.5 47.7
LCFD-C-B32 64.7 48.4 30.6 71.0 53.7
4.6 Further Discussion
During this entire discussion, all experiments are based on the strong version of LCFD, i.e., LCFD-C-B32. For brevity, we remove the suffix ’-C-B32’.
4.6.1 Analysis for Causal Factors Discovery
[…]
Figure 5: Accuracy changing of pseudo-label as training iteration going on.
Figure 5: Accuracy changing of pseudo-label as training iteration going on.
Figure 6: Stacked area chart of accuracy as Gaussian blur intensity enhancing.
Analysis of discovering process. In our LCFD framework, the latent external causal elements $S_{e}$ are captured and transformed into the observed pseudo-label. Thereby, the pseudo-label quality should be an indicator reflecting the discovery process. Here, we select the pseudo-label-based methods SHOT, COWA, GKD and CLIP’s zero shot (termed CLIP) as evaluation references. The comparison results are depicted in Fig. 6. Due to conducting $S_{e}$ discovery with the aid of CLIP, the pseudo-label accuracy of LCFD, i.e., LCFD-PL, is much better than previous methods as well as CLIP-PL. On the other hand, LCFD’s curve is on the top of LCFD-PL with a significant gap during the entire training process. This phenomenon shows that capturing internal causal elements $S_{i}$ is effective.
Causality evaluation of discovered causal factors. Ideally, when a model is designed to establish underlying causal relationships accurately, the model will exhibit immunity/ robustness to noise [27]. To evaluate the captured causal factors, we conduct a robustness analysis to noise. Specifically, we observe the accuracy change, whilst gradually increasing the kernel size of Gaussian blur (from 8 to 20) on the input target images. The comparison methods are CLIP, SHOT and LCFD-w-CLIP, which is a variation of LCFD where the pseudo-labels adopt the zero-shot results of CLIP directly.
As shown in Fig. 6, the performance of LCFD-w-CLIP, SHOT, CLIP deteriorates to varying degrees, as noise increases. In contrast, LCFD almost maintains the performance on all intensities. Specifically, compared with the accuracy at kernel size equals 8, LCFD-w-CLIP, SHOT, CLIP respectively decrease by 5.0%, 9.1% and 7.9% when kernel size equals 20, i.e., max noise intensity, whilst the decrease of LCFD is only 0.3%. The lower performance loss of LCFD implies the effectiveness of capturing causal factors by LCFD. In particular, together LCFD with CLIP, LCFD can work with impressive noise robustness as CLIP’s zero shot performance is sharply weakening. This result implies that LCFD’s performance is not due to CLIP’s excellent zero-shot capability, but deriving from the discovery of causal factors.
Also, comparing LCFD with LCFD-w-CLIP, the gap between them gradually broadens, implying capturing the external elements $S_{e}$ enables LCFD to gain immunity to noise. To gain a better understanding, we visualize the prediction distribution of LCFD-PL and CLIP at zero and max noise intensity, using the t-SNE tool. Fig. 7 displays the visualization results.
Figure 7: Prediction visualization of LCFD-PL, CLIP when Gaussian blur kernel sizes are 0 (N0) and 20 (N20), respectively. Their corresponding classification accuracy (%) is bracketed.
[…]
When Gaussian blur kernel size is 0 (the first row), the clustering shape of LCFD-PL-N0 and CLIP-PL-N0 have a similar pattern: Samples with ambiguous semantics are mainly distributed in the middle area. As the kernel size reaches 20, CLIP-N20 continues the previous distribution pattern that ambiguous samples spread to a much wider central zone. Thereby, the noise drowns out the accurate category information provided by generated pseudo labels. As a result, the LCFD-w-CLIP’s performance deteriorates.
On the contrary, despite the accuracy being almost the same as CLIP-N20 (31%), LCFD-PL-N20 evolves a new pattern: While these ambiguous samples move to the periphery of the distribution, the remaining samples still cluster in the middle area with a shape similar to the LCFD-PL-N0 case. That is, as the noise enhancing, LCFD-PL-N20 maintains the distribution information. Correspondingly, with LCFD-PL-N20 guidance, LCFD achieves robustness to noise.
4.6.2 Model Analysis
Feature distribution visualization. Feature distribution is an intuitive evaluation means for classification. Here, we select five comparisons, including the source model (termed Source), CLIP, SHOT, TPDS and Oracle (trained on domain Cl with the real labels). Fig. 8 shows the logit visualization results in 3D Density chart. From Source to LCFD, the clustering effect gradually strengthens, and LCFD is the most similar to Oracle.
Ablation study. In the proposed framework, we discover the causal factors $S=\{S_{e},S_{i}\}$ in a separate way. We isolate the effect of this strategy by three variation LCFD methods. (i) LCFD-w-$S_{i}$: We only take $\cal{L}_{\rm{UN}}$ in Eq. (14) as the objective for model training, such that we test the effect of $S_{i}$ capturing. (ii) LCFD-w-$S_{e}$: We only perform the pseudo-label supervised learning, regulated by $\cal{L}_{\rm{PL}}$, where the pseudo-label is generated from $S_{e}$ discovery regulated by $\cal{L}_{\rm{EC}}$. Thus, the effect of $S_{e}$ capturing is tested. (iii) LCFD-w-CLIP: We directly adopt the zero-shot results of CLIP as the pseudo-labels regulating the $S_{i}$ discovery.
TABLE VIII: Ablation study results (%). CS, OS, PS and GS mean the closed-set, open-set, partial-set and Generalied SFDA settings, respectively.
Method Office-Home IN-K
CS OS PS GS Avg. SF-OODG
Source 59.2 46.6 62.8 73.1 60.4 30.2
LCFD-w-$S_{i}$ 70.3 71.5 70.1 75.9 72.6 36.9
LCFD-w-$S_{e}$ 81.6 80.7 84.1 84.3 82.2 44.2
LCFD-w-CLIP 80.9 78.3 83.9 82.5 80.6 44.2
LCFD 83.3 83.4 85.8 84.5 83.7 48.4
TABLE IX: Scalability study (%) of IC instantiation.
Loss Method Closed-set SF-OODG
Office-Home VisDA IN-K
$\cal{L}_{\rm{UN}}$ LCFD+ENT 82.5 89.4 46.8
LCFD+NRC 83.0 87.9 47.5
LCFD+TPDS 83.5 90.2 45.0
[…]
LCFD+${\rm{KL}}$ 83.5 89.1 49.1
LCFD 83.3 89.2 48.4
Tab. VIII reports the accuracy of these methods above in the four SFDA settings and SF-OODG. Specifically, when LCFD-w-$S_{i}$ or LCFD-w-$S_{e}$ works, the result outforms Source, but significantly lower than the full version LCFD. The results indicate that individual capturing of $S_{e}$ or $S_{i}$ is effective, and they jointly contribute to the final performance. Also, the comparison of LCFD-w-CLIP and LCFD confirm the effect of $\cal{L}_{\rm{EC}}$-based prompt learning for $S_{e}$ discovery.
Scalability study of IC instantiation. The objective $\cal{L}_{\rm{IC}}$ in Eq. (14) regulating $S_{i}$ discovery is a salable component with multiple implementations. To verify its scalability, this part respectively gives three other instantiations for the two regulators in $\cal{L}_{\rm{IC}}$. For $\cal{L}_{\rm{UN}}$, (i) the alone entropy regulating without category balance constraint (LCFD+ENT), (ii) data local structure-based regulating (LCFD+NRC) and (iii) domain shift control-based regulating (LCFD+TPDS) are selected as alternatives. For $\cal{L}_{\rm{PL}}$, the popular 1-Norm ($||\cdot||_{1}$), 2-Norm ($||\cdot||_{2}$) and KL divergence are employed.
As the results listed in Tab. IX, in the closed-set setting, LCFD+TPDS obtain the best results, and other methods have a gap of up to 2%. This indicates that these typical regulators do not substantially impact model performance in the closed-set setting. In the SF-OODG setting, besides LCFD+${\rm{KL}}$, all methods have a certain accuracy decrease compared with LCFD. For LCFD+ENT, removing category balance regulating leads to a decrease. The decrease of LCFD+NRC and LCFD+TPDS is due to the method not specifically considering the SF-OODG setting. Regarding the $\cal{L}_{\rm{PL}}$ group, the results indicate that probability distribution approximation (${\rm{KL}}$) outshines the features closing ($||\cdot||_{1}$, $||\cdot||_{2}$) in terms of maintaining the model performance caused by the external causal elements.
Effect of prompt initialization. In the LCFD method, we adopt prompt learning to encode the external causal factors $S_{e}$. In this part, we give four kinds of initialization strategies to figure out the effect of prompt initiation. As listed in Tab. X, except for conventional (i) constant strategy (row 1$\sim$2), (ii) sentence strategy (row 3$\sim$4) and (iii) phrase strategy (row 5$\sim$7), we introduce an interesting strategy of sentence initialization with uncertainty (row 8$\sim$9), considering the lacking of accurate supervision in SFDA.
Tab. X gives the comparison results of nine templates in the closed-set setting and SF-OODG. It can be observed that these templates do not lead to significant performance variance. These results indicate that the LCFD model is insensitive to prompt selection. Meanwhile, comparing the constant strategy and other strategies, the initialization with semantics is a better option, which also meets our expectations.
TABLE X: Prompt initialization study (%).
# Initialization template Closed-set SF-OODG
Office-Home VisDA IN-K
1 ’X [CLASS].’ (#X=4) 82.6 89.1 48.4
2 ’X [CLASS].’ (#X=16) 81.5 88.5 46.1
3 ’There is a [CLASS].’ 83.6 89.2 48.6
4 ’This is a photo of a [CLS].’ 83.7 89.0 48.8
5 ’This is maybe a photo of a [CLS].’ 83.0 89.1 48.8
6 ’This is almost a photo of a [CLS].’ 83.7 89.2 48.4
7 ’This is definitely a photo of a [CLS].’ 83.5 88.9 48.1
8 ’a picture of a [CLS].’ 83.5 88.8 48.6
9 ’a photo of a [CLS].’ 83.3 89.2 48.4
[…]
5 Conclusion
This paper considers a more challenging Unified SFDA problem that thoroughly addresses all specific scenarios in a unified manner. Further, we propose a LCFD approach from a causality view, conceptually different from prior statistical dependence-based scheme. Specifically, we propose a self-supervised information bottleneck to capture the external ones and further transform them into pseudo-labels to regulate the successive IC process for discovering internal ones. According to our theoretical result, we instantiate the proposed information bottleneck based on the network implementation for variational mutual information. Extensive experiments on four mainstream SFDA settings and SF-OODG validate the superiority of our LCFD.
Acknowledgments
This work is partially supported by the German Research Foundation and National Natural Science Foundation of China (NSFC) in project Crossmodal Learning under contract Sonderforschungsbereich Transregio 169, the Hamburg Landesforschungsförderungsprojekt Cross, NSFC (61773083); Horizon2020 RISE project STEP2DYNA (691154); NSFC (62206168, 62276048).